# Sustainable and Precision Agriculture with the Internet of Everything (IoE)

### v2

@@ Abstract
The accelerated pace of global population growth underscores the crucial role of the agricultural sector in mitigating food scarcity, as well as in supporting livelihoods through employment opportunities and bolstering national economies. This sector faces several critical challenges, including resource depletion, socioeconomic issues, gaps in technology and innovation, and the impact of climate change. The introduction of mechanization has significantly transformed agriculture by enhancing sustainability and increasing the productivity of crops. Recently, traditional farming methods have been supplemented with advanced technologies such as artificial intelligence (AI), robotics, biotechnology, blockchain, and Internet of Things (IoT), steering the industry towards precision agriculture. The convergence of these advanced technologies has facilitated the automation of various tasks such as water management, crop monitoring, disease management, and harvesting.
Advancements in IoT technology have been instrumental in addressing the longstanding challenges faced by agriculture, particularly those arising from resource limitations and climate change. The concept of Internet of Everything (IoE) has gained traction due to its holistic approach towards integrating various IoT specializations, called “IoXs” where X referring to a specific domain. This includes areas like the Internet of Sensors (IoS), Internet of Vehicles (IoV), Internet of Energy (IoEn), Internet of Space Things (IoST), Industrial Internet of Things (IIoT), and Internet of Drones (IoD). This paper explores the potential of the Internet of Everything (IoE) in revolutionizing agricultural systems. The focus is on assessing the impact of cutting-edge and novel technologies, such as 6G, molecular communication (MC), Internet of Nano Things (IoNT), Internet of Bio-Nano Things (IoBNT), Internet of Fungus, and designer phages, in significantly improving agricultural yield, efficiency, and productivity. Additionally, the potential of these technologies is evaluated in terms of their applicability, associated challenges, and future research directions within the realm of precision agriculture.
@@ I Introduction
Agriculture is crucial for human sustenance and equally serves as a cornerstone of a nation’s economic structure. It catalyzes agribusiness ventures that stimulate job creation and increase foreign exchange revenues through exports [1]. This pivotal role of agriculture in modern economies has its roots in historical evolution. The first agricultural revolution occurring around 10,000 B.C, played a key role in enabling a significant increase in human population [2].The British Revolution spanning the 17th–19th centuries, witnessed a dramatic upswing in agricultural production in Britain [3], setting the stage for Industrial Revolution. The Green Revolution of the mid 20th century led to a marked increase in agricultural yields due to a series of research, development, and technology transfer initiatives [4]. The emergence
@@
of Agriculture 4.0 represents the integration of modern information and communication technologies (ICT) with classic agricultural methods [5]. This progression has laid the groundwork for the upcoming phase of agricultural evolution, emphasizing the merging of various technologies, especially wireless ones. Fig. 1 illustrates this progression through key periods of agricultural revolution.
Agricultural technology has progressed in tandem with agricultural practices; the adoption of wireless systems in this sector marked a turning point in agricultural management and optimization strategies. Wireless communication is now among the most rapidly evolving, dynamic and vibrant technological areas in recent times [6]. Future advancements focus on achieving extensive area coverage, enhanced capacity, high data rates, reduced latency, superior quality of service, and minimal energy consumption as key objectives [7]. According to current literature [8] present wireless technologies are being promoted with the capability to provide wireless speeds of up to ten gigabits per second (10 Gbps). Even with such advancements, the unequal implementation of technology remains a substantial obstacle, with rural communities at risk of falling even more behind in terms of digital access [9]. This necessitates focused attention on three areas: firstly, increasing awareness about the use of communication technologies in agriculture; secondly, establishing a very robust communication network; and thirdly, fostering skill development.
Over recent years, notable progress in wireless communication systems has primarily stemmed from innovations in their designs. This advancement is a result of implementation of various modulation techniques, sophisticated coding methods, utilization of multiple-input multiple-output (MIMO) technologies and other cutting edge strategies [10]. However, challenges such as constrained bandwidth, latency related delays, interference from external sources , and weak security protocols are still affecting both data flow and user experience in multitude of applications. Despite these challenges, significant advancements in wireless technology are being made. The introduction of Wi-Fi 6 (802.11ax) technology is addressing congestion problems in scenarios that require a significant amount of bandwidth, while also improving efficiency and speed [11]. At the same time, the deployment of 5G networks offers extremely-low latency and high speeds, making it possible for realisation of applications that span from smart agriculture and autonomous vehicles to smart urban development [12]. Additionally, the development of mesh networks [13] is contributing to the scalability and reliability of wireless systems.
Building on these technological improvements, wireless technologies have enabled internet access to billions of people around the world, thereby effectively integrating a vast global population into the contemporary digital economy. These technologies are vital in numerous sectors like banking, agriculture, transportation, and healthcare, transforming communication and business operations in the digital age [14]. In agriculture, specifically, wireless technology stands at the forefront of advancements. It significantly enhances the management of water resources for irrigation [15], improves pest management strategies [16], facilitates balanced fertilization [17] , and enables efficient greenhouse monitoring [18], among other applications. These diverse applications demonstrate the significant role of wireless technology, which extends beyond merely offering global connectivity to actively promoting notable developments in fields like agriculture.
The Internet of Things (IoT) sector has emerged as a formidable catalyst in the advancement of technology, reshaping how interconnected objects are used for identification, sensing, data collection and communicating [19]. By integrating real-time data processing and analytics with advanced connectivity, it promises substantial improvements in functionality and efficiency across various fields, especially in agriculture. According to [20], the basic concept of an IoT system is centered on machine to machine (M2M) data transfer driven by state-of-the-art technologies like wireless sensor networks (WSN) and radio frequency identification (RFID). The framework of IoT is structured around three principal visions: internet-based, object-based, and semantic-based [21]. This system employs sensing devices equipped with intelligent algorithms for effective decision-making. A comprehensive bibliometric study of IoT has been presented in [22], which explores various challenges associated with IoT implementation and also proposes viable solutions. This transformation towards a more efficient and accessible world is effectively depicted in Fig. 2. This figure includes a block diagram and details the various applications areas of IoT technology.
Agriculture is a primary beneficiary of the IoT revolution, with IoT introducing numerous innovative opportunities for modernizing this sector. The integration of IoT is recognized as a vital strategy for enhancing agriculture,
@@
a field that demands constant monitoring and management. Currently, farmers and growers spend approximately 70% of their time tracking and understanding crop conditions [23], a task that IoT can streamline, thereby freeing up time for actual agricultural labor. Key applications of IoT in agriculture include precision agriculture, livestock surveillance, and greenhouse management, each falling into unique monitoring domains. The implementation of WSN, along with assorted IoT sensors and devices, facilitates the collection of critical data by farmers and researchers, significantly optimizing agricultural practices [24]. Thus, integration of IoT with traditional farming techniques ultimately leads to a new phase of agricultural innovation, creating higher standards for efficiency, sustainability, and yield maximization.
The Internet of Everything (IoE) extends the IoT concept to an integrated network of vast array of computing devices, sensors, along with both inanimate and living entities, people, data and processes. This IoE network, encompassing both tangible & intangible entities, creates an interconnected ecosystem capable to function as a seamless and self-regulating network. The innovative IoE concept envisions the entire universe as interconnected, ranging from molecules to plants, animals, people, buildings, vehicle, and drones, thus representing a holistic perspective of universal connectivity. This holistic view is set to present myriad of possibilities, like engaging at the molecular level with plants and animals and boosting the efficiency and productivity of smart agriculture.
The IoE vision seeks to connect pervasive electronic devices to the Internet, analyze the extensive data they generate, and provide intelligent applications to benefit human society. To attain this vision, IoE aims to realise three vital expectations:
  • Scalability: Establish a network with pervasive coverage.
  • Intelligence: Create a global computing infrastructure for smart decision-making.
  • Diversity: Support a variety of applications across different domains/ sectors.
Despite the advancements and presence of enabling technologies aimed at fulfilling these key objectives, IoE still faces a host of challenges. These challenges range from the high costs associated with hardware to limitations in network infrastructure, from the scarcity of available spectrum to computational constraints, as well as interoperability and pervasive security concerns. In [25] the concept of IoE Framework has been introduced highlighting the specialization of IoT into specific application domains, or IoXs, with ‘X’ denoting various sectors or domains such as the Internet of Agricultural Things, Internet of Energy or the Internet of Vehicles etc. This framework fosters collaborations among various IoXs.
Additionally, the study discusses the prospective applications of the IoE model and the various challenges associated with its implementation. In smart farming, IoE’s transformative potential lies in its integration with IoXs and emerging technologies, enhancing agricultural practices through improved resource management, precise monitoring, and data-driven decision-making, thereby fostering a more sustainable and productive agricultural future.
While technology plays a cruical role in agriculture, it is also important to recognize and address the various ongoing challenges and limitations that persist within this sector. Modern ICT has transformed agriculture but the digital technology dividends are not automatic. Particularly, rural areas often lag in adoption of these technologies due to poor infrastructure, a lack of expertise and awareness. Furthermore, the adoption of these technologies also raises certain ethical, legal and social issues including ownership & privacy of data, accountability, and potential health & environmental impacts.
For small-scale farmers, the initial investment and maintenance costs associated with these modern technologies such as wireless sensors, drones, robots etc., can be excessively high. To alleviate this, the government could offer affordable loans to help these farmers manage the upfront costs. Enabling technologies, like IoT and artificial intelligence (AI), have significantly transformed traditional agricultural practices [26]. However, the extensive power requirements of wireless sensors and central servers present a significant challenge. Solutions might include the deployment of energy harvesting techniques and a shift towards more autonomous edge devices or sensors capable of conducting the necessary computations independently. The incorporation of IoT and similar technologies in smart agriculture also introduces cyber-security risks, potentially compromising farmers’ privacy and the security of critical crop and equipment data. Blockchain technology could be a viable option to secure the data while it is being transmitted from wireless sensors to the cloud servers. Additionally, designing robust and fault tolerant IoT hardware and equipment that needs to withstand harsh and extreme environmental conditions is a crucial concern [27],[28],[29]. Interoperability among IoT devices and hardware is an additional critical issue, especially as the number of these devices and platforms grows exponentially. In this regard, development of universal standards and protocols is urgently needed, with regulatory bodies advocating for standardization to ensure compliance.
Another significant challenge is the potential of considerable losses in terms of production and time due to incorrect diagnosis of plant diseases by sensors. This necessitates the development of precise image processing techniques coupled with machine learning algorithms capable of accurately identifying the various infections of plants throughout their life-cycle, thereby mitigating potential losses and enhancing agricultural efficiency. Embracing these modern technologies while acknowledging and addressing their associated challenges and limitations will bring revolutionary advancements in the agriculture sector.
@@
The paper begins by introducing the concept of critical role of agriculture in human sustenance and economic development, tracing its evolution through various technological eras. It then highlights the importance of technologies, especially IoT, in agricultural domain. Additionally, the transformative impact of IoE on the agriculture sector was examined. The remainder of the paper is organized as follows: section II presents an overview of the Internet of Agricultural Things (IoAT), detailing various application areas, challenges and future research opportunities. Section III focuses on the application of nanotechnology in enhancing smart agriculture. Section IV investigates harnessing phages to improve animal and plant health. Section V explores the impact of Internet of Fungus on contemporary agriculture. Section VI & VII delve into the Internet of Energy Harvesting Things and the integration of the Internet of Vehicles, Drones, and Space Things in achieving sustainable and efficient farming practices. Sections VIII to X discuss the transformative potential of 6G, blockchain technology and machine learning in advancing agricultural practices. The paper concludes with Section XI, summarizing the key findings and insights.
II Internet of Agricultural Things (IoAT)

### v3

@@ Abstract
Agriculture faces critical challenges from population growth, resource scarcity, and climate change, driving a shift toward advanced, technology-integrated farming. Mechanization has transformed agriculture, enhancing sustainability and crop productivity. Now, technologies like artificial intelligence (AI), robotics, biotechnology, blockchain, and Internet of Things (IoT) are advancing precision agriculture. The concept of Internet of Everything (IoE) has gained traction due to its holistic approach towards integrating various IoT specializations, called “IoXs” where X referring to a specific domain. This paper explores the transformative role of the IoE in agriculture, expanding beyond traditional IoT applications to integrate niche subdomains like molecular communication (MC), the Internet of Nano Things (IoNT), Internet of Bio-Nano Things (IoBNT), designer phages and Internet of Fungus (IoF). Our study provides a detailed review of how these IoE subdomains, in conjunction with 6G, blockchain, and machine learning (ML), can enhance precision farming in areas like crop monitoring, resource management, and disease control. Unlike prior IoT-centric reviews, this work uniquely focuses on IoE’s potential to advance agriculture at molecular and biological scales, achieving more precise resource utilization and resilience. Key contributions include an exploration of these technologies’ applicability, associated challenges, and recommendations for future research directions within precision agriculture.
@@ I Introduction
Agriculture is crucial for human sustenance and a foundational pillar of national economies, stimulating agribusiness, job creation, and generating foreign exchange through exports. Its importance is historically established, with milestones such as the first agricultural revolution around 10,000 B.C. [1] and the British Revolution in the 17th –19th centuries [2], which led to significant productivity increases during the latter period. Later, the mid-20th century Green Revolution further boosted agricultural yields through research, development, and technology transfer [3]. Today, Agriculture 4.0 integrates information and communication technologies (ICT) with traditional agricultural practices, setting the stage for a new phase of agricultural evolution focused on sustainability and efficiency [4]. Fig. 1 illustrates this agricultural evolution, highlighting advancements from crop rotation and high-yield crops to modern precision technologies.
@@
Technological advancements in wireless communication have significantly transformed agriculture by enabling extensive coverage, increased capacity, high data rates, and lower latency. Emerging technologies, such as Wi-Fi 7 and 5G, address challenges related to bandwidth, efficiency and speed, supporting applications in smart agriculture and urban development [5]. However, digital inequality in rural areas remains a significant barrier to adoption, highlighting the need for improved infrastructure, network resilience, and technical skill development [6].
The Internet of Things (IoT) has emerged as a formidable catalyst in the advancement of technology, allowing for real-time data processing, analytics, and advanced connectivity, which are particularly valuable in agriculture for applications like precision farming, livestock monitoring, and greenhouse management [7], [8]. By creating a network of interconnected devices and data-driven decision-making processes, IoT advances the goals of Agriculture 4.0. This transformation towards a more efficient and accessible world is effectively depicted in Fig. 2. This figure includes a block diagram and details the various applications areas of IoT technology.
@@
Expanding on IoT, the Internet of Everything (IoE) connects diverse entities-such as sensors, data, and living organisms-into a unified network. This approach enables scalability, intelligence, and cross-domain functionality with potential applications like molecular-level interactions with plants and animals. The IoE Framework, as introduced in [9], specializes IoT into specific domains, or IoXs, such as the Internet of Agricultural Things (IoAT), the Internet of Energy (IoEn), or the Internet of Vehicles (IoV), etc., fostering collaboration across domains. In smart farming, IoE’s potential lies in integrating IoXs and emerging technologies to optimize resource management, monitoring, and decision-making, supporting a more sustainable and productive agricultural future.
This study aims to investigate the transformative potential of the IoE in enhancing precision agriculture. While numerous survey and review articles have explored IoT’s role in smart and precision agriculture, they primarily focus on conventional IoT systems and wireless communication. To the best of our knowledge, this study is the first to thoroughly investigate the integration of both unconventional IoE subdomains- such as molecular communication (MC), the Internet of Bio-Nano Things (IoBNT), designer phages, & the Internet of Fungus (IoF)- and emerging technologies like 6G and machine learning (ML) etc. within agricultural applications. This work identifies the unique challenges posed by each technology and presents a forward-looking perspective on potential research directions to harness these innovations for sustainable and efficient agriculture.
@@
The paper begins by introducing the concept of critical role of agriculture in human sustenance and economic development, tracing its evolution through various technological eras. It then highlights the importance of technologies, especially IoT, in agricultural domain. Additionally, the transformative impact of IoE on the agriculture sector was examined. The remainder of the paper is organized as follows : Section II outlines the research methodology, encompassing the framework, motivation, and research questions guiding the study of IoE applications in agriculture. Section III presents an overview of IoAT, detailing various application areas, challenges and future research opportunities. Section IV focuses on the application of nanotechnology in enhancing smart agriculture. Section V investigates harnessing phages to improve animal and plant health. Section VI explores the impact of IoF on contemporary agriculture. Section VII & VIII delve into the Internet of Energy Harvesting Things and the integration of the Internet of Vehicles, Drones, and Space Things in achieving sustainable and efficient farming practices. Sections IX through XI discuss the transformative potential of 6G, blockchain technology and ML in advancing agricultural practices. Section XII provides discussion and analysis of critical factors for successful IoE adoption in agriculture. Finally, the paper concludes with Section XIII, summarizing the key findings and insights. To assist readers, the organization of this article is illustrated in Fig. 3, with acronyms listed in Table I.
TABLE I: List of Acronyms
Acronym Description
IoAT Internet of Agricultural Things
ICT Information and Communication Technologies
IoE Internet of Everything
AI Artificial Intelligence
IoT Internet of Things
IoBNT Internet of Bio-Nano Things
IoNT Internet of Nano Things
IoD Internet of Drones
IoV Internet of Vehicles
IoST Internet of Space Things
ML Machine Learning
6G Sixth Generation (Wireless Technology)
IIoT Industrial Internet of Things
IoEn Internet of Energy
PCA Principal Component Analysis
SVD Singular Value Decomposition
MDP Markov Decision Process
DQN Deep Q Network
SVM Support Vector Machine
MC Molecular Communication
V2V Vehicle-to-Vehicle Communication
V2X Vehicle-to-Everything Communication
VANET Vehicle Ad-hoc Networks
IRS Intelligent Reflecting Surface
NFV Network Function Virtualization
CSA Climate-Smart Agriculture
THz Terahertz
RIS Reconfigurable Intelligent Surface
UAV Unmanned Aerial Vehicle
GPS Global Positioning System
LWP Leaf Water Potential
PAR Photosynthetically Active Radiation
VRT Variable Rate Technology
NP Nanoparticles
IoS Internet of Sensors
VOCs Volatile Organic Compounds
RFID Radio-frequency Identification
Figure 3: Organization / Structure of this article
II Research Methodology
A well-structured research methodology is key to ensuring the rigor and reliability of any study. It provides a systematic framework that guides the research process, facilitating precise data collection, in-depth analysis, and insightful interpretation to effectively address specific research questions.This section outlines the formulation of research motivation, the development of relevant & specific research questions, a search string strategy, screening of search results, and the steps of data extraction and evaluation, as depicted in Fig.4.
Figure 4: Fundamental stages of systematic literature review
II-A Motivation
The motivation for conducting this study is grounded in several compelling reasons, which are enumerated below:
  1. 1.
    The rapid advancement of technology is transforming various sectors, including agriculture. This shift calls for moving beyond traditional agricultural practices to enhance productivity and efficiency. Consequently, it is essential to provide an overview of the application areas where ICT can drive smart and precision agriculture.
  2. 2.
    The integration of IoT as a core component in realizing IoAT introduces multiple challenges that demand effective management or complete resolution to optimize the efficacy and productivity of agriculture.
  3. 3.
    Existing literature reviews and survey articles typically focus on conventional IoT and wireless communication advancements in the agricultural domain. However, there is an urgent need to uniquely explore niche, uncharted technologies and their potential transformative impacts on agriculture.
The main objective of this review paper is to addresses the critical gap in understanding how these various technologies in the IoE domain can advance precision agricultural.
II-B Research Questions
Defining specific research questions (RQs) is essential when conducting a literature review, as it provides a clear direction and structured roadmap to facilitate the selection of relevant articles and data for analysis. For this study, the following RQs have been established to guide the research process effectively.
  • RQ1: What is the importance of ICT in precision agriculture?
  • RQ2: What are the application areas of IoT in smart agriculture?
  • RQ3: What are the current and emerging trends of IoT in smart agriculture?
  • RQ4: What are the key challenges specific to IoT in this sector?
  • RQ5: What novel IoXs can be employed in precision agriculture?
  • RQ6: What are the limitations of these IoXs?
  • RQ7: What are the potential research direction of these IoXs?
  • RQ8: What are the critical factors influencing the successful adoption and integration of IoE in agriculture?
II-C Search Strings
To gather the most relevant and recent studies aligned with the RQs and objectives of this paper, specific search terms and structured search strategies were employed. Precise keywords were essential in narrowing the search and focusing on relevant literature. Table II lists the search engines, digital libraries and databases, while Table III presents key words/ string searches used in this research. It is worth noting that only articles published in English were considered.
II-D Related Works
Over the past decade, substantial research has advanced the field of smart agriculture, particularly through the integration of IoT technologies for enhanced productivity and sustainability. Table IV presents a curated selection of key studies from the last five years, chosen based on their relevance to the convergence of IoT with emerging IoE subdomains, as explored in this work. Criteria for selection include the studies’ focus on technical challenges and solutions related to IoT-driven precision agriculture, their in-depth analysis of innovative applications in smart farming, and the extent to which they identify future research directions. This focused approach enables a deeper examination of the transformative potential and practical applications of IoE technologies within agriculture, addressing both opportunities and challenges unique to the field.
TABLE II: Search Engines & Digital Libraries
Source URL
IEEE https://ieeexplore.ieee.org/
FRONTIERS https://www.frontiersin.org/
ELSEVIER https://www.elsevier.com/
SPRINGER https://www.springer.com/gp
MDPI https://www.mdpi.com/
NATURE https://www.nature.com/
Scopus https://www.scopus.com/
Google Scholar https://scholar.google.com/
Research Gate https://www.researchgate.net/
TABLE III: List of Search Strings
Databases Search String/ Keywords
Sources in Table II (Internet of Everything OR IoE) AND (Precision Agri OR Smart Agri) OR (Internet of Things OR IoT) AND (Precision Agri OR Smart Agri) OR (IoT agricultural survey) OR (IoE agricultural survey) OR (IoE based agriculture applications OR precision agriculture applications OR Machine learning models OR ML for smart agriculture applications) OR (Blockchain OR BC smart agriculture) AND (IoE OR Internet of Everything)
TABLE IV: Related Works - Precision / Smart Agriculture
Reference Key Enabling Technology Research Objective Application Domain
[10] 2024 Computer vision, AI, ML, imaging. Analyzes computer vision techniques across crop life cycles for precision farming, focusing on automation and decision-making. Crop monitoring, disease detection, yield prediction.
[11] 2023 IoT, blockchain, AI, ML, robotics Provides an overview of smart agriculture fundamentals, tech enablers, and potential future developments. Crop management, food security, sustainable practices.
[12] 2023 IoT, wireless communication, sensors, mechanization Examines IoT and sensor integration to improve automation and control in agricultural practices. Crop health monitoring, irrigation management, yield improvement.
[13] 2022 IoT, robotics, WSN, AI. Systematically reviews Agriculture 4.0 technologies, adoption rates, and key barriers to digitization. Crop management, in-field monitoring, automation.
[14] 2022 IoT, big data, AI, blockchain, gene sequencing. Reviews IoT and AI-driven data solutions for enhancing monitoring, food safety, and traceability in agri-food systems. Crop monitoring, food quality, supply chain traceability.
[15] 2022 IoT, AI, UAVs, WSNs Provides a critical review of IoT and AI advancements, deployment challenges, and future trends in smart agriculture. Smart irrigation, pest detection, crop monitoring.
[16] 2022 IoT, UAVs, remote sensing, data analytics Discusses IoT frameworks, components, security issues, and future trends for automating agricultural management. Automated irrigation, soil monitoring, pest detection
[17] 2022 IoT, WSN, agbots, drones, AI. Reviews IoT-enabled sensor technologies and their integration with advanced computing for effective smart agriculture. Soil monitoring, pest detection, climate adaptation.
[18] 2022 ML, IoT, robotics, drones, sensors. Investigates ML applications to enhance automation, crop health monitoring, and resource efficiency in agriculture. Precision farming, disease detection, irrigation, yield prediction.
[19] 2022 WSN, IoT, automation Explores IoT applications and challenges in monitoring and control for smart agriculture. Climate monitoring, soil fertility, pest detection, irrigation.
[20] 2021 IoT, AI, blockchain, cybersecurity, edge computing. Examines smart agriculture’s development modes, enabling technologies, and security/privacy challenges. Precision farming, facility agriculture, CSA.
[21] 2021 IoT, UAVs, SDN, NFV, cloud computing Provides a comprehensive overview of emerging IoT technologies and their applications for advancing smart agriculture. Smart monitoring, disease management, supply chain, water management.
[22] 2021 IoT, CSA practices, data analytics. Analyzes CSA adoption on small-scale farms, focusing on productivity, resilience, and GHG reduction. Small-scale farming, climate resilience, productivity gains.
[23] 2021 IoT, cloud computing, ML, AI, UAVs. Examines smart farming technologies, including IoT and AI, to address agriculture’s productivity and climate impact challenges. Crop and animal production, post-harvest management, climate resilience.
[24] 2021 IoT, AI, drones, imaging (multispectral, hyperspectral) To review and synthesize advanced control and monitoring strategies in smart agriculture for automation, crop health, and precision farming. Stress monitoring, irrigation management, disease detection, yield prediction.
[25] 2020 AI, ML, WSN, imaging, robotics Investigates AI-driven cognition tools to address crop health, pest control, and resource optimization in agriculture. Precision farming, disease detection, crop phenotyping.
[26] 2020 IoT, robotics, AI, big data, VRT Reviews advancements in data management for smart farming, emphasizing automation and sustainability. Crop data management, precision farming, sustainable practices.
[27] 2019 UAVs, AI, big data, IoT, robotics. Reviews UAV technology for agricultural tasks, emphasizing platform types, control systems, and future applications. Crop mapping, pesticide spraying, crop monitoring.
[28] 2019 AI, IoT, ML, wireless comms To assess AI-driven automation techniques addressing issues like crop diseases, irrigation, and pest control in smart farming. Disease detection, irrigation optimization, yield enhancement.
[29] 2019 Edge computing, fog computing, IoT Examines edge computing’s role in overcoming connectivity challenges for scalable smart agriculture applications. Livestock health, crop production, farm security.
This Review IoAT, MC, IoNT, IoBNT, designer phages, IoF, ML, 6G, IoD, IoV, IoEn, Blockchain Investigates the integration of IoE in agriculture, focusing on novel subdomains and highlighting challenges and opportunities for enhancing precision agriculture. Precision resource management, pest control, nutrient exchange, disease detection.
II-E Contributions
This study provides a pioneering perspective on IoE-enabled precision agriculture by investigating the integration of unconventional IoE technologies in agricultural systems. It delivers a forward-looking analysis of precision farming, emphasizing advanced technologies that push the boundaries of agricultural innovation. By identifying key research directions, this study positions itself as a foundational reference for advancing IoE in agriculture, distinguishing it from previous research focused solely on IoT. The primary contributions of this article are outlined as follows.
  1. 1.
    Identification of IoT Application Areas in Smart Agriculture: This study identifies the primary areas where IoT technologies can be applied, such as crop monitoring, disease control, livestock management, and soil nutrient management. By systematically examining these areas, the study reveals the diversity of IoT in addressing diverse agricultural challenges and enhancing operational efficiency.
  2. 2.
    Exploration of Novel IoE Subdomains in Agriculture: Our study is the first to systematically investigate how unconventional IoE subdomains can be applied within precision agriculture. Unlike traditional IoT, these technologies operate at molecular and biological scales, allowing for advanced precision in resource management, pest control, and disease detection. Key innovations compared to existing literature review / survey papers are as follows.
    • Focus on Molecular and Bio-Nano Communication in agriculture.
    • Introduction of designer phages for sustainable disease control.
    • Importance of the IoF for enhanced nutrient and information exchange.
  3. 3.
    Identification of Technical Challenges and Research Opportunities: This study identifies the core technical and operational challenges associated with implementing IoE at scale in agriculture, emphasizing sustainable solutions. It also presents key future research directions to guide the effective integration of these technologies into agricultural systems.
III Internet of Agricultural Things